Question 3 Sunland, Inc. manufactures ergonomically designed computer furniture. Sunland uses a job order costing system. On November 30, the Work in Process Inventory consisted of the following jobs: Job No. Item Units Accumulated Cost CC723 Computer caddy 21,000 $930,000 CH291 Chair 15,000 439,000 PS812 Printer stand 27,000 253,000 $1,622,000 On November 30, Sunland’s Raw Materials Inventory account totaled $697,000, and its Finished Goods Inventory totaled $3,469,400. Sunland applies manufacturing overhead on the basis of machine hours. The company’s manufacturing overhead budget for the year totaled $4,700,000, and the company planned to use 940,000 machine hours during the year. Through the first eleven months of the year, the company used a total of 845,000 machine hours, total manufacturing overhead amounted to $4,294,500, and Cost of Goods Sold was $4,274,500. Sunland purchased $8,750,250 in raw materials in December and incurred the following costs for jobs in process that month: Job No. Materials Issued Machine Hours Direct Labor Hours Direct Labor Cost CC723 $185,000 14,000 13,700 $128,400 CH291 $18,800 4,900 3,800 $44,200 PS812 $248,000 20,600 17,300 $228,500 DS444 $259,000 17,000 13,500 $150,000 The following jobs were completed in December and transferred to the Finished Goods Inventory: Job No. Item Units CC723 Computer caddy 21,000 CH291 Chair 15,000 DS444 Desk 6,000 Assume that Sunland sold 16,000 computer caddies, 12,700 chairs, and 4,400 desks in December. Collapse question part

(a) Sunland incurred $395,000 in manufacturing overhead in December. Given that expenditure, was manufacturing overhead under- or overapplied for the year? By how much? Overhead by $
Answer #1

Budgeted manufacturing overhead for the year = $4,700,000

Budgeted machine-hours for the year = 940,000

Predetermined manufacturing overhead rate = $4,700,000/940,000 = $5 per machine-hour

---------------------------------------------------------------------------------------------------------------------------

Machine-hours used during the first 11 months = 845,000

Machine-hours used in December = 14,000 + 4,900 + 20,600 + 17,000 = 56,500

Therefore,

Total machine hours used during the year = 845,000 + 56,500 = 901,500

---------------------------------------------------------------------------------------------------------------------------

Manufacturing overhead applied during the year

= Total machine-hours used x Predetermined overhead rate

= 901,500 x $5

= $4,507,500

----------------------------------------------------------------------------------------------------------------------------

Manufacturing overhead cost incurred during the year

= Manufacturing overhead cost incurred till November 30 + Manufacturing overhead cost incurred in December

= $4,294,500 + $395,000

= $4,689,500

-----------------------------------------------------------------------------------------------------------------------------

From the above calculations it can be seen that the amount of manufacturing overhead applied during the year is less than the amount of manufacturing overhead cost incurred.

Therefore, manufacturing overhead is underapplied for the year.

Manufacturing overhead is underapplied by $182,000 ($4,689,500 - $4,507,500).